OpenStudy (anonymous):

The point (2, 2) is a solution to the system of equations 4y – 4x = 0 and y - 6 = -2x. TRUE OR FALSE

OpenStudy (anonymous):

true because: when you replace y and x by (2,2) it equal 0 in both side

OpenStudy (anonymous):

was that helpful?

OpenStudy (anonymous):

4y-4x=0 4(2)-4(2)=0 8-8=0 0=0

OpenStudy (anonymous):

and do that for the rest